OpenStudy (anonymous):

HELP!!!!!!!! What is the average value of y=e^x between x=e and x=2e?

OpenStudy (amistre64):

the slope of the line between 2 points ...

OpenStudy (amistre64):

\[\frac{y(e)-y(2e)}{e-2e}\]

OpenStudy (anonymous):

These are the options that I have to choose from: I. e II. e^(2e) - e^(e) III. e^(2e-1) - e^(e-1) IV. e^(2e-1) V. e^(2)

OpenStudy (anonymous):

I got that the answer is V. e^(2), but I think that is not correct.

OpenStudy (amistre64):

\[\Large \frac{e^e-e^{2e}}{e-2e}\] \[\Large \frac{e(e^{e-1}-e^{2e-1})}{-e}\] \[\Large e^{2e-1}-e^{e-1}\]

OpenStudy (anonymous):

Thank you so much! I really appreciate it...!

OpenStudy (amistre64):

i think i broke a synapse with that one :)

OpenStudy (anonymous):

Hahahaha! :-)

OpenStudy (anonymous):

 I. f(x) is continuous and differentiable everywhere II. f(x) is continuous but not differentiable at x = e III. f(x) is differentiable but not continuous at x = e IV. f(x) is neither continuous nor differentiable V. f(x) is undefined at x = e